OpenStudy (anonymous):

A board game has a spinner with ten equal sections, numbered 1 through ten. In order to win the game, Mae needs to spin at least a 6 on each of her last two turns. What's the probability of this happening? one percent ten percent twenty-five percent thirty-six percent

OpenStudy (anonymous):

at least a 6 means 6,7,8,9,10 so five out of ten or one half for each try since she has to do it twice, that would be one half times one half or \[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\]

OpenStudy (anonymous):

so 25%?

OpenStudy (anonymous):

yes

OpenStudy (mrnood):

@satellite73 It is discouraged by the code of conduct for this site to give direct answers. The OP learns nothing from having someone else give them the answer. Maybe you could've stopped before doing the sum above - your explanation would give the OP a chance to try to do the sum for himself....

OpenStudy (anonymous):

@MrNood -_- he is being helpful and he did not even give out a direct answer

OpenStudy (anonymous):

what are u gonna do ban him ???? xD he's a mod